OpenStudy (jordan123321):

What is the value of the expression x + (–y) when x = –3.8 and y = 1.7? A. –5.5 B. –2.1 C. 2.1 D. 5.5

OpenStudy (anonymous):

substitute the given values into the equation

OpenStudy (anonymous):

−3.8−1.7 −3.8−1.7 =−5.5

OpenStudy (anonymous):

\[x+(-y)=-3.8-1.7\]

OpenStudy (anonymous):

\[=?\]

OpenStudy (anonymous):

solve it @jordan123321

OpenStudy (jordan123321):

i cant tht is why i asked

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

−3.8−1.7 −3.8−1.7 =−5.5 @jordan123321

OpenStudy (anonymous):

\[-3.8-1.7=?\]

OpenStudy (jordan123321):

what is tht^^

OpenStudy (anonymous):

−3.8−1.7 −3.8−1.7 =−5.5 @jordan123321 The answer^

OpenStudy (anonymous):

\[x+(-y)=x-y\]

OpenStudy (anonymous):

substitute x=-3.8 and y=1.7 into the equation

OpenStudy (anonymous):

\[=x-y\]

OpenStudy (anonymous):

\[=-3.8-1.7\]

OpenStudy (anonymous):

\[=?\]

OpenStudy (anonymous):

@jordan123321